OpenStudy (anonymous):

Suppose you just bought a car. Your payments are $185 a month, your insurance is $115 a month, and it costs you $0.37 per mile to operate. Write an equation that represents the total monthly cost of your vehicle to drive it x miles per month.

OpenStudy (experimentx):

200+0.37*x

OpenStudy (anonymous):

aNSWER CHOICES y = 0.37x + 115 y = 0.37x + 300 y = 185x + 115 y = 0.37x + 185

OpenStudy (experimentx):

Oops!! sorry, I made a mistake ... the first one y = 0.37*x+300

OpenStudy (anonymous):

No problem sweetie (: